4. Consider the one-period binomial model, and let V1 S1- That is, the derivative security pays the stock price at time t-1. Find the time t = 0 no-arbitrage price of the derivative, Vo
Answer #1

There are multiple ways to solve this problem.

Method 1:

The value of the derivative at t=1 is V1 = S1

That means i have to create a replicating portfolio today that will result into a value of S1 at the end of time period 1. The replicating portfolio should be nothing but 1 number of stock today. This portfolio of 1 number of stock will assume the value of S1 at the end of period 1.

So, V0 = value of the derivative today = value of the replicating portfolio today = value of 1 number of stock today = S0

Hence, V0 = S0

-------------------

Method 2:

The symbols have usual meanings.

Let's revisit the assumptions and the results of 1 period binomial model where a stock with current price of S0 can move to Su = S0 x u in up state and Sd = S0 x d in down state.

Assume continuous compounding risk free interest rate r.

Risk neutral probability of up state at the end of 1 period: Pu = (er - d) / (u - d) and

Risk neutral probability of down state at the end of 1 period: Pd = (u - er) / (u - d)

Expected value of the stock at the end of 1 period, S1 = Pu x Su + Pd x Sd

(e - d) (u - d) (d-e) (u - d)

Hence the expected value of the derivative at the end of 1 period = V1 = S1 = S0 x er

Hence value of the derivative today = V0 = PV of V1 = V1.e-r = S0 x er x e-r = S0